OpenStudy (anonymous):

The set {-1, 0, 1} is not closed under which operation?

Parth (parthkohli):

It's not closed under division, because \(1\div 0\) or \(-1 \div 0\) maps to \(\text{undefined}\).

Parth (parthkohli):

If you try other operations, you'd map to an element present in the set.

Parth (parthkohli):

\[0 \times 1 = 0 \\ 1 + (-1) = 0\\ -1 - 0 = -1 \\ \text{etc.} \]

Parth (parthkohli):

@Riza Do you get it?

OpenStudy (anonymous):

here is the things i can choose as an answer. multiplication, division, subtraction, all of the above

Parth (parthkohli):

Just go through what I told you.

Parth (parthkohli):

Closure property just states that you map to an element present in the set. Oh c'mon! You have done this in basic arithmetic too!

OpenStudy (anonymous):

i think it multiplication

Parth (parthkohli):

Why multiplication? Multiply any two numbers in the set, and you get a number that is present in the set.

Parth (parthkohli):

\[1 \times 0 = 0 \]\[1 \times -1 = -1 \]\[-1 \times 0 = 0 \]

OpenStudy (anonymous):

The set {0,1} is not closed under subtraction because we cannot subtract 1 from 0 in that set

Parth (parthkohli):

lol, the question mentions the set {-1,0,1}.

OpenStudy (anonymous):

bleh.

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

Now I think you have left with only one option @Riza

OpenStudy (anonymous):

its not multiplication division or subtraction

OpenStudy (anonymous):

But you have guessed for Multiplication and Subtraction only..

Parth (parthkohli):

\[-1 \div 1 = -1 \]\[-1 \div 0 = \text{underfined} \]Note how undefined is not given in the set.

OpenStudy (anonymous):

Can you tell what is the value of: \[\frac{1}{0} = ??\]